# Decentralized CBF-based Safety Filters for Collision Avoidance of Cooperative Missile Systems with Input Constraints

### v1

@@ III Problem Formulation
While proportional navigation guidance (PNG) provides efficient interception trajectories, its standard formulation does not explicitly consider interactions among multiple effectors. Consequently, agents pursuing different targets may inadvertently converge into close proximity, creating the risk of mutual collision. This limitation motivates the integration of a real-time safety filter that modifies the nominal guidance command into a safe acceleration command $\mathbf{a}_{i}$, thereby ensuring both feasibility and satisfaction of safety constraints.
@@ IV Decentralized Robust CBF-based QP Formulation for Collision Avoidance
To anticipate potential future collisions, we compute the zero-effort miss distance (ZEM), which quantifies the predicted minimum separation between two agents assuming they maintain their current velocities. This quantity is derived from the relative motion between agents $i$ and $j$, under the assumption of constant velocity propagation. The relative position at time $t=t_{0}+T$ evolves as
@@
$\min_{\mathbf{a}_{i}\in\mathbb{R}^{3}}\;\;\lVert\mathbf{a}_{i}(t)-\mathbf{a}_{\text{nom},i}(t)\rVert_{W}^{2}$
@@ V Criticality-Based Prioritization under Conflicting Constraints
$\min_{\mathbf{a}_{i}\in\mathbb{R}^{3},\{\delta_{ij}\geq 0\}}\quad\|\mathbf{a}_{i}-\mathbf{a}_{\text{nom},i}\|^{2}+\sum_{j\in\mathscr{N}_{i}}w_{ij}\,\delta_{ij}$ (39)
@@ VI Simulation Results
To evaluate the proposed decentralized RCBF-based safety filter, we consider a three-on-three interception scenario with proportional navigation guidance (PNG) serving as the nominal law. Each effector is assigned to a unique target, resulting in one-to-one pairings throughout the engagement. The evaluation compares two cases: (i) a baseline scenario using PNG-only guidance and (ii) a scenario where PNG is combined with the proposed safety filter.

### v2

@@ III Problem Formulation
While PNG provides efficient interception trajectories, its standard formulation does not explicitly consider interactions among multiple effectors. Consequently, agents pursuing different targets may inadvertently converge into close proximity, creating the risk of mutual collision. This limitation motivates the integration of a real-time safety filter that modifies the nominal guidance command into a safe acceleration command $\mathbf{a}_{i}$, thereby ensuring both feasibility and satisfaction of safety constraints.
@@ IV Decentralized Robust CBF-based QP Formulation for Collision Avoidance
To anticipate potential future collisions, we compute the ZEM distance, which quantifies the predicted minimum separation between two agents assuming they maintain their current velocities. This quantity is derived from the relative motion between agents $i$ and $j$, under the assumption of constant velocity propagation. The relative position at time $t=t_{0}+T$ evolves as
@@
$\min_{\mathbf{a}_{i}\in\mathbb{R}^{3}}\;\;\lVert\mathbf{a}_{i}-\mathbf{a}_{\text{nom},i}\rVert_{W}^{2}$
@@ V Criticality-Based Prioritization under Conflicting Constraints
$\min_{\mathbf{a}_{i}\in\mathbb{R}^{3},\{\delta_{ij}\geq 0\}}\quad\|\mathbf{a}_{i}-\mathbf{a}_{\text{nom},i}\|_{W}^{2}+\sum_{j\in\mathscr{N}_{i}}w_{ij}\,\delta_{ij}$ (39)
@@ VI Simulation Results
To evaluate the proposed decentralized RCBF-based safety filter, we consider a three-on-three interception scenario with PNG serving as the nominal law. Each effector is assigned to a unique target, resulting in one-to-one pairings throughout the engagement. The evaluation compares two cases: (i) a baseline scenario using PNG-only guidance and (ii) a scenario where PNG is combined with the proposed safety filter.